At 1 atm of pressure water boils at 100°C, if the sample was placed under 2 atm of pressure, what would be the temperature? (This would be like a pressure cooker).


Law: ______________________


Answer:____________________


On hot days, you may have noticed that potato chip bags seem to “inflate”, even though they have not been opened. If I have a 250mL bag at a temperature of 19 °C, and I leave it in my car which has a temperature of 60 °C, what will the new volume of the bag be?


Law: ______________________


Answer:____________________


In a certain experiment a sample of helium in a vacuum system was compressed at 25 °C from a volume of 200.0 mL to a volume of 0.240 mL where its pressure was found to be 30.0 mm Hg. What was the original pressure of the helium? 


Law: ______________________


Answer:____________________




1
Expert's answer
2021-04-29T07:22:27-0400

1. Gaylussac's law

P1/T1 = P2/T2

1/373 = 2/T2

0.002681 = 2/T2

0.002681 T2 = 2

745.99K - 273

= 472.99°C

2. Ideal gas law

V1/T1 = V2/T2

250/292 = V2/333

0.8562 × 333 = V2

= 285.10 ml

3. Gaylussac's law

P1V1 = P2V2

P1 × 200 = 240 × 30

200P1 = 720

= 3.6 atm